uxx - (sech^4x)uyy = 0
Reduce into canonical form and solve it
1
Expert's answer
2020-12-15T07:08:40-0500

"u_{xx}-sech^4xu_{yy}=0"


"A=1, B=0, C=-sech^4x"

"\Delta=B^2-4AC=4sech^4x>0"

This is hyperbolic PDE.


The characteristic polynomial of the PDE:

"A(\frac{dy}{dx})^2-B(\frac{dy}{dx})+C=0"


"\frac{dy}{dx}=\frac{B\pm\sqrt{B^2-4AC}}{2A}=\pm sech^2x"

"y=tanhx+k_1"

"y=-tanhx+k_2"


"\xi=y-tanhx"

"\eta=y+tanhx"


We have:

"a=0, c=0"


Canonical form:

"\omega_{\xi\eta}=0"


"\omega(\xi,\eta)=f(\xi)+g(\eta)"


The general solution:

"u=f(y-tanhx)+g(y+tanhx)"